OpenStudy (anonymous):

Consider the line that passes through the point (-3, 2) and has a slope of 3. Part 1: Write the equation of this line using point-slope form. (2 points) Part 2: Using your equation from part 1, rewrite this equation in slope-intercept form. Make sure to show all of your work. (2 points) Part 3: Using your equation from part 2, rewrite this equation in standard form. Make sure to show all of your work. (2 points)

OpenStudy (anonymous):

Wait what?

OpenStudy (anonymous):

part 1 y - y1 = m( x - x1) put y1 = 2 and x1 = -3 and m = 3

OpenStudy (anonymous):